# Zihan Zhang Princeton

### v4

@@ 2 Problem formulation
  • Hypothesis class. Suppose we are interested in a hypothesis class $\mathcal{H}$, comprising a set of candidate functions from the instance space $\mathcal{X}$ to the label space $\mathcal{Y}$. Overloading the notation, we use $h_{\pi}$ to represent a randomized hypothesis associated with a probability distribution $\pi\in\Delta(\mathcal{H})$, meaning that a hypothesis $h$ from $\mathcal{H}$ is randomly selected according to distribution $\pi$. When it comes to a VC class, we assume that the hypotheses in $\mathcal{H}$ are binary-valued, and that the VC dimension (vapnik1994measuring) of $\mathcal{H}$ is
@@
  • Loss function. Suppose we are given a loss function :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$, so that $\ell(h,(x,y))$ quantifies the risk of using hypothesis $h\in\mathcal{H}$ to make prediction on a datapoint $(x,y)\in\mathcal{X}\times\mathcal{Y}$ (i.e., predicting $y$ based on $x$). One example is the 0-1 loss function $\ell(h,(x,y))=\mathds{1}\{h(x)\neq y\}$, which is often used to measure the misclassification error.
@@
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h_{\pi}\sim\pi}\big{[}\ell\big{(}h_{\pi},(x,y)\big{)}\big{]}\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon.$ (7)
@@ 3 Algorithm
       /* recompute $\widehat{w}^{t}$ draw new samples for $\mathcal{S}$ only if $w^{t}$ changes sufficiently. */
@@ 9                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
Following the game dynamics proposed in previous works (awasthi2023sample; haghtalab2022demand), our algorithm alternates between computing the most favorable hypothesis (performed by the learner) and estimating the most challenging mixture of data distributions (performed by the adversary), with the aid of no-regret learning algorithms (roughgarden2016twenty; shalev2012online). More specifically, in each round $t$, our algorithm performs the following two steps:
@@
    Given hypothesis $h^{t}$, we compute an updated weight vector $w^{t+1}\in\Delta(k)$ — and hence an updated mixed distribution $\mathcal{D}^{(t+1)}=\sum_{i\in[k]}w_{i}^{t+1}\mathcal{D}_{i}$. The weight updates are carried out using the celebrated Hedge algorithm (freund1997decision) designed for online adversarial learning, [^5] Note that the Hedge algorithm is closely related to Exponentiated Gradient Descent, Multiplicative Weights Update, Online Mirror Descent, etc (arora2012multiplicative; shalev2012online; hazan2022introduction). in an attempt to achieve low regret even when the loss vectors are adversarially chosen. More precisely, we run
@@
In words, the min-player and the max-player in our algorithm follow the best-response dynamics and no-regret dynamics, respectively (note that the Hedge algorithm is known to be a no-regret algorithm (roughgarden2016twenty)). At the end of the algorithm, we output a randomized hypothesis $h^{\sf final}$ that is uniformly distributed over the hypothesis iterates $\{h^{t}\}_{1\leq t\leq T}$ over all $T$ rounds, following common practice in online adversarial learning.
While the above paradigm has been adopted in past works (awasthi2023sample; haghtalab2022demand), the resulting sample complexity depends heavily upon how data samples are collected and utilized throughout the learning process. For instance, awasthi2023sample — which also alternates between best response and no-regret algorithm — draws fresh data at each step of every round, in order to ensure reliable estimation of the loss function of interest through elementary concentration inequalities. This strategy, however, becomes wasteful over time, constituting one of the main sources of its sample sub-optimality.
@@
  • Sample reuse in Step (a). In stark contrast to awasthi2023sample that draws new samples for estimating each $h^{t}$, we propose to reuse all samples collected in Step (a) up to the $t$-th round to assist in computing $h^{t}$. As will be made precise in lines 1-1 of Algorithm 1, we shall maintain a growing dataset $\mathcal{S}$ for conducting Step (a) throughout, ensuring that there are $n_{i}^{t}$ samples drawn from distribution $\mathcal{D}_{i}$ in the $t$-th round. These datapoints are employed to construct an empirical loss estimator $\widehat{L}^{t}(h,w^{t})$ for each $h\in\mathcal{H}$ in each round $t$, with the aim of achieving uniform convergence $|\widehat{L}^{t}(h,w^{t})-L(h,w^{t})|\leq O(\varepsilon)$ over all $h\in\mathcal{H}$. More detailed explanations are provided in Section 4.1.
@@ 4.1 Towards sample reuse: uniform concentration and a key quantity
Armed with this result, we can extend it to accommodate all $h\in\mathcal{H}$ through the union bound. For a VC class with $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d$, the celebrated Sauer–Shelah lemma (wainwright2019high, Proposition 4.18) tells us that the set of hypotheses can be effectively compressed into a subset with cardinality no larger than $\exp\big{(}\widetilde{O}(d)\big{)}$. Taking the union bound then yields
@@
where the equivalence arises from von Neumann’s minimax theorem (v1928theorie). Consider the following algorithm (cf. Algorithm 2) tailored to this minimax problem, assuming perfect knowledge about the loss vectors. [^6] Note that in Algorithm 1, we can only estimate the loss vector using the collected samples. Additional efforts are needed to reduce the variability (see line 1 in Algorithm 1).
@@ 4.2.1 First attempt: bounding the number of distributions with $\max_{1\leq t\leq T}w^{t}_{i}=\Omega(1)$
To do so, note that some standard “continuity”-type argument tells us that: for a sufficiently small stepsize $\eta$, one can find, for each $i\in\mathcal{W}_{\mathsf{large}}$, a time interval $[s_{i},e_{i}]\subseteq[0,T]$ obeying
@@
Let us briefly discuss the high-level proof ideas. Following standard analysis for the Hedge algorithm (e.g., shalev2012online; lattimore2020bandit), we can often obtain (under certain mild conditions)
@@
Combine this relation with basic properties about the KL divergence (see, e.g., Lemmas 10 and 11 in Appendix A) and the choice $\eta=\varepsilon/10$ to obtain
@@ 4.2.2 More general cases: issues and solutions
Nevertheless, simply repeating the arguments in Section 4.2.1 fails to deliver the desirable bound (23) on $|\mathcal{W}(p)|$ when $p$ is small. Briefly speaking, for each $i\in\mathcal{W}(p)$, let $[s_{i},e_{i}]$ represent a time interval (akin to (19)) such that
@@
  • Shared intervals. For the interval $[s_{i},e_{i}]$ associated with $i$ as defined above, if there exist other indices sharing the same interval $[s_{i},e_{i}]$ (in the sense that relations analogous to (24) are satisfied for other indices), then it is plausible to improve the bound. For instance, if a set $\mathcal{M}_{i}\subseteq[k]$ of indices share the same interval $[s_{i},e_{i}]$, then one can follow the heuristic argument in (25) to obtain
@@
      for each $n\in[N]$, all indices in $\mathcal{V}_{n}$ share the same interval $[s_{n},e_{n}]$ (defined analogously as (24));
@@
    $|\mathcal{W}_{j}|=\sum_{n=1}^{N}|\mathcal{V}_{n}|\leq\sum_{n=1}^{N}O\bigg{(}\frac{(e_{n}-s_{n})\varepsilon^{2}}{p}\bigg{)}\leq O\bigg{(}\frac{T\varepsilon^{2}}{p}\bigg{)}=\widetilde{O}\bigg{(}\frac{1}{p}\bigg{)}.$ (28)
In light of the above discussion, it is helpful to (a) merge those indices that share similar intervals, and (b) identify disjoint intervals whose associated indices can cover a good fraction of $\mathcal{W}_{j}$.
@@
In general, however, it is common for two segments to be overlapping (see Figure 2 in Appendix C.5), which precludes us from directly invoking our aforementioned observation about “disjoint intervals.” To address this issue, our strategy is to extract out shared sub-segments [^7] A sub-segment refers to a sub-interval of a segment, as illustrated in Figure 6. of (a subset of) these segments in a meticulous manner. Encouragingly, it is possible to find such sub-segments that taken collectively cover a good fraction (i.e., $\frac{1}{\mathrm{poly}\log(k,T)}$) of all segments, meaning that we do not have to discard too many segments. The construction is built upon careful analysis of these segments, and will be elucidated in Appendix C.
@@ 5 Analysis for VC classes (proof of Theorem 1)
With probability at least $41-\delta/$ 4,
@@
Suppose that lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability exceeding $41-\delta/$ 4, the hypothesis $h^{\mathsf{final}}$ output by Algorithm 1 is $\varepsilon$-optimal in the sense that
$\max_{1\leq i\leq k}L(h^{\mathsf{final}},\mathrm{e}^{\mathsf{basis}}_{i})\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon.$ (30)
@@
Taking Lemma 1 and Lemma 2 together, one can readily see that Algorithm 1 returns an $\varepsilon$-optimal randomized hypothesis $h^{\mathsf{final}}$ with probability at least $21-\delta/$ 2. The next step then lies in bounding the total number of samples that has been collected in Algorithm 1. Towards this end, recall that $\overline{w}^{T}_{i}=\max_{1\leq t\leq T}w_{i}^{t}$ for each $i\in[k]$. Recognizing that $\widehat{w}^{t}_{i}\leq\overline{w}^{t}_{i}$ for each $t\in[T]$ and $i\in[k]$, we can bound the total sample size by
(sample $(\text{sample size})\qquad T_{1}\sum_{i=1}^{k}\widehat{w}_{i}^{T}+k+T\bigg{(}k\sum_{i=1}^{k}\overline{w}_{i}^{T}+k\bigg{)}$ $\leq\left(T_{1}\|\overline{w}^{T}\|_{1}+kT\|\overline{w}^{T}\|_{1}\right)+k(T+1)$
@@
Assume that lines 1-1 in Algorithm 1 are replaced with some oracle which returns a hypothesis $h^{t}$ satisfies that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability at least $41-\delta/$ 4, the quantity $\|\overline{w}^{T}\|_{1}$ is bounded above by
@@
It is noteworthy that the proof of Lemma 3 is the most technically challenging part of the analysis; we postpone this proof to Appendix B.3.
@@
as claimed in Theorem 1. It remains to prove the above key lemmas, which we postpone to Appendix B.
6 Necessity of randomization
Given that the best-known sample complexities prior to our work were derived for algorithms that either output randomized hypotheses or invoke majority votes, awasthi2023sample raised the question about how the sample complexity is impacted if only deterministic (or “proper”) hypotheses are permitted as the output of the learning algorithms. As it turns out, the restriction to deterministic hypotheses substantially worsens the sample efficiency, as revealed by the following theorem.

### v5

@@ 2 Problem formulation
  • Hypothesis class. Suppose we are interested in a hypothesis class $\mathcal{H}$, comprising a set of candidate functions from the instance space $\mathcal{X}$ to the label space $\mathcal{Y}$. Overloading the notation, we use $h_{\pi}$ to represent a randomized hypothesis associated with a probability distribution $\pi\in\Delta(\mathcal{H})$, meaning that a hypothesis $h$ from $\mathcal{H}$ is randomly selected according to distribution $\pi$. When it comes to a VC class, we assume that the hypotheses in $\mathcal{H}$ are binary-valued, and that the VC dimension (Vapnik et al.,, 1994) of $\mathcal{H}$ is
@@
  • Loss function. Suppose we are given a loss function $\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$, so that $\ell(h,(x,y))$ quantifies the risk of using hypothesis $h\in\mathcal{H}$ to make prediction on a datapoint $(x,y)\in\mathcal{X}\times\mathcal{Y}$ (i.e., predicting $y$ based on $x$). One example is the 0-1 loss function $\ell(h,(x,y))=\mathds{1}\{h(x)\neq y\}$, which is often used to measure the misclassification error.
@@
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h\sim\pi}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon.$ (7)
@@ 3 Algorithm
  /* recompute $\widehat{w}^{t}$ & draw new samples for $\mathcal{S}$ only if $w^{t}$ changes sufficiently. */
@@ 9        $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
Following the game dynamics proposed in previous works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), our algorithm alternates between computing the most favorable hypothesis (performed by the learner) and estimating the most challenging mixture of data distributions (performed by the adversary), with the aid of no-regret learning algorithms (Roughgarden,, 2016; Shalev-Shwartz,, 2012). More specifically, in each round $t$, our algorithm performs the following two steps:
@@
    Given hypothesis $h^{t}$, we compute an updated weight vector $w^{t+1}\in\Delta(k)$ — and hence an updated mixed distribution $\mathcal{D}^{(t+1)}=\sum_{i\in[k]}w_{i}^{t+1}\mathcal{D}_{i}$. The weight updates are carried out using the celebrated Hedge algorithm (Freund and Schapire,, 1997) designed for online adversarial learning, [^5] Note that the Hedge algorithm is closely related to Exponentiated Gradient Descent, Multiplicative Weights Update, Online Mirror Descent, etc (Arora et al.,, 2012; Shalev-Shwartz,, 2012; Hazan,, 2022). in an attempt to achieve low regret even when the loss vectors are adversarially chosen. More precisely, we run
@@
In words, the min-player and the max-player in our algorithm follow the best-response dynamics and no-regret dynamics, respectively (note that the Hedge algorithm is known to be a no-regret algorithm (Roughgarden,, 2016)). At the end of the algorithm, we output a randomized hypothesis $h^{\sf final}$ that is uniformly distributed over the hypothesis iterates $\{h^{t}\}_{1\leq t\leq T}$ over all $T$ rounds, following common practice in online adversarial learning.
While the above paradigm has been adopted in past works (Awasthi et al.,, 2023; Haghtalab et al.,, 2022), the resulting sample complexity depends heavily upon how data samples are collected and utilized throughout the learning process. For instance, Awasthi et al., (2023, Algorithm 1) — which also alternates between best response and no-regret algorithm — draws fresh data at each step of every round, in order to ensure reliable estimation of the loss function of interest through elementary concentration inequalities. This strategy, however, becomes wasteful over time, constituting one of the main sources of its sample sub-optimality.
@@
  • Sample reuse in Step (a). In stark contrast to Awasthi et al., (2023, Algorithm 1) that draws new samples for estimating each $h^{t}$, we propose to reuse all samples collected in Step (a) up to the $t$-th round to assist in computing $h^{t}$. As will be made precise in lines 1-1 of Algorithm 1, we shall maintain a growing dataset $\mathcal{S}$ for conducting Step (a) throughout, ensuring that there are $n_{i}^{t}$ samples drawn from distribution $\mathcal{D}_{i}$ in the $t$-th round. These datapoints are employed to construct an empirical loss estimator $\widehat{L}^{t}(h,w^{t})$ for each $h\in\mathcal{H}$ in each round $t$, with the aim of achieving uniform convergence $|\widehat{L}^{t}(h,w^{t})-L(h,w^{t})|\leq O(\varepsilon)$ over all $h\in\mathcal{H}$. More detailed explanations are provided in Section 4.1.
@@ 4.1 Towards sample reuse: uniform concentration and a key quantity
Armed with this result, we can extend it to accommodate all $h\in\mathcal{H}$ through the union bound. For a VC class with $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})=d$, the celebrated Sauer–Shelah lemma (Wainwright,, 2019, Proposition 4.18) tells us that the set of hypotheses can be effectively compressed into a subset with cardinality no larger than $\exp\big{(}\widetilde{O}(d)\big{)}$. Taking the union bound then yields
@@
where the equivalence arises from von Neumann’s minimax theorem (v. Neumann,, 1928). Consider the following algorithm (cf. Algorithm 2) tailored to this minimax problem, assuming perfect knowledge about the loss vectors. [^6] Note that in Algorithm 1, we can only estimate the loss vector using the collected samples. Additional efforts are needed to reduce the variability (see line 1 in Algorithm 1).
@@ 4.2.1 First attempt: bounding the number of distributions with $\max_{1\leq t\leq T}w^{t}_{i}=\Omega(1)$
To do so, note that some standard “continuity”-type argument tells us that: for a sufficiently small stepsize $\eta$, one can find, for each $i\in\mathcal{W}_{\mathsf{large}}$, a time interval $s_{i},e_{i}]\subseteq[0,T]$ obeying
@@
Let us briefly discuss the high-level proof ideas. Following standard analysis for the Hedge algorithm (e.g., Shalev-Shwartz, (2012); Lattimore and Szepesvári, (2020)), we can often obtain (under certain mild conditions)
@@
Combine this relation with basic properties about the KL divergence (see, e.g., Lemmas 15 and 16 in Appendix A) and the choice $\eta=\varepsilon/10$ to obtain
@@ 4.2.2 More general cases: issues and solutions
Nevertheless, simply repeating the arguments in Section 4.2.1 fails to deliver the desirable bound (23) on $|\mathcal{W}(p)|$ when $p$ is small. Briefly speaking, for each $i\in\mathcal{W}(p)$, let $s_{i},e_{i}]$ represent a time interval (akin to (19)) such that
@@
  • Shared intervals. For the interval $s_{i},e_{i}]$ associated with $i$ as defined above, if there exist other indices sharing the same interval $s_{i},e_{i}]$ (in the sense that relations analogous to (24) are satisfied for other indices), then it is plausible to improve the bound. For instance, if a set $\mathcal{M}_{i}\subseteq[k]$ of indices share the same interval $s_{i},e_{i}]$, then one can follow the heuristic argument in (25) to obtain
@@
      for each $n\in[N]$, all indices in $\mathcal{V}_{n}$ share the same interval $s_{n},e_{n}]$ (defined analogously as (24));
@@
    $|\mathcal{W}(p)|=\sum_{n=1}^{N}|\mathcal{V}_{n}|\leq\sum_{n=1}^{N}O\bigg{(}\frac{(e_{n}-s_{n})\varepsilon^{2}}{p}\bigg{)}\leq O\bigg{(}\frac{T\varepsilon^{2}}{p}\bigg{)}=\widetilde{O}\bigg{(}\frac{1}{p}\bigg{)}.$ (28)
In light of the above discussion, it is helpful to (a) merge those indices that share similar intervals, and (b) identify disjoint intervals whose associated indices can cover a good fraction of $\mathcal{W}(p)$.
@@
In general, however, it is common for two segments to be overlapping (see Figure 2 in Appendix B.6), which precludes us from directly invoking our aforementioned observation about “disjoint intervals.” To address this issue, our strategy is to extract out shared sub-segments [^7] A sub-segment refers to a sub-interval of a segment, as illustrated in Figure 6. of (a subset of) these segments in a meticulous manner. Encouragingly, it is possible to find such sub-segments that taken collectively cover a good fraction (i.e., $\frac{1}{\mathrm{poly}\log(k,T)}$) of all segments, meaning that we do not have to discard too many segments. The construction is built upon careful analysis of these segments, and will be elucidated in Appendix 5.3.
@@ 5.1 Key lemmas underlying the proof
With probability at least $1-\delta/4$,
@@
Suppose that lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability exceeding $1-\delta/4$, the hypothesis $h^{\mathsf{final}}$ output by Algorithm 1 is $\varepsilon$-optimal in the sense that
$\max_{1\leq i\leq k}L(h^{\mathsf{final}},\mathrm{e}^{\mathsf{basis}}_{i})\leq\min_{\pi\in\Delta(\mathcal{H})}\max_{1\leq i\leq k}L(h_{\pi},\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon.$ (30)
@@
Taking Lemma 1 and Lemma 2 together, one can readily see that Algorithm 1 returns an $\varepsilon$-optimal randomized hypothesis $h^{\mathsf{final}}$ with probability at least $1-\delta/2$. The next step then lies in bounding the total number of samples that has been collected in Algorithm 1. Towards this end, recall that $\overline{w}^{T}_{i}=\max_{1\leq t\leq T}w_{i}^{t}$ for each $i\in[k]$. Recognizing that $\widehat{w}^{t}_{i}\leq\overline{w}^{t}_{i}$ for each $t\in[T]$ and $i\in[k]$, we can bound the total sample size by
$(\text{sample size})\qquad T_{1}\sum_{i=1}^{k}\widehat{w}_{i}^{T}+k+T\bigg{(}k\sum_{i=1}^{k}\overline{w}_{i}^{T}+k\bigg{)}$ $\leq\left(T_{1}\|\overline{w}^{T}\|_{1}+kT\|\overline{w}^{T}\|_{1}\right)+k(T+1)$
@@
Assume that lines 1-1 in Algorithm 1 are replaced with some oracle which returns a hypothesis $h^{t}$ satisfies that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability at least $1-\delta/4$, the quantity $\|\overline{w}^{T}\|_{1}$ is bounded above by
@@
It is noteworthy that the proof of Lemma 3 is the most technically challenging part of the analysis, and we shall present this proof in Section 5.2.
@@
as claimed in Theorem 1.
5.2 Proof of Lemma 3 (controlling the Hedge trajectory)
As alluded to previously, the proof of Lemma 3 forms the most technically challenging part of our analysis. Let us begin by introducing several convenient notation. Set $\delta^{\prime}=\delta/(32T^{4}k^{2})$, and let
Let us define
$\overline{\mathcal{W}}$ $\coloneqq[k]/\cup_{j}\mathcal{W}_{j}.$ (33b)
Lemma 4.
holds all $1\leq j\leq\overline{j}$, with $\overline{j}$ defined in (32).
Importantly, this lemma tells us that, with probability at least $1-8T^{4}k^{2}\delta^{\prime}=1-\delta/4$, one has
Noteworthily, proving Lemma 4 is the most challenging part of our analysis, and we dedicate the next subsection (Section 5.3) to the proof of Lemma 4.
5.3 Proof of Lemma 4 (bounding $|\mathcal{W}_{j}|$ for each $j$)
The proof relies heavily on the concepts of “segments” introduced in Section 4.2. Throughout this section, we shall take $\delta^{\prime}=\delta/(32T^{4}k^{2})$, and focus on any $j$ obeying
$1\leq j\leq\log_{2}(k)-2.$ (35)
Before continuing, we find it helpful to underscore a high-level idea: if $|\mathcal{W}_{j}|$ is large, then there exist many disjoint segments, thereby requiring the total length $T$ to be large enough in order to contain these segments. The key steps to construct such disjoint segments of interest are as follows:
  1. 1.
    Construct a suitable segment for each $i\in\mathcal{W}_{j}$ (see Lemma 5);
  2. 2.
  3. 3.
5.3.1 Step 1: showing existence of a segment for each distribution in $\mathcal{W}_{j}$
Lemma 5.
Proof.
$\tau=\max\big{\{}t\mid t\leq e_{i},w^{t}_{i}\leq 2^{-(j+2)}\big{\}},$
5.3.2 Step 2: constructing disjoint segments with good coverage
Lemma 6.
  1. $\mathrm{(i)}$
  2. $\mathrm{(ii)}$
  3. $\mathrm{(iii)}$
Proof.
See Appendix B.3. ∎
Lemma 7.
Recall the definition of $\mathcal{W}_{j}$ in (33a). Then there exists a group of subsets $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ of $\mathcal{W}_{j}$ satisfying the following properties:
  1. $\mathrm{(i)}$
  2. $\mathrm{(ii)}$
  3. $\mathrm{(iii)}$
    • $\mathrm{(a)}$
    • $\mathrm{(b)}$
    • $\mathrm{(c)}$
Proof.
See Appendix B.4. ∎
Remark 3.
The endpoints $\{\widehat{s}_{n}\}$ and $\{\widehat{e}_{n}\}$ mentioned above are carefully constructed in the proof. The quantities $\{g_{n}\}$ are only implicitly defined here, since their precise values (except the property $g_{n}\geq 1$) are not needed in the subsequent analysis; the interested reader can find more precise details about $\{g_{n}\}$ in the proof.
In brief, Lemma 7 identifies a collection of subsets $\{\mathcal{V}_{j}^{n}\}_{n=1}^{N}$ of $\mathcal{W}_{j}$ enjoying the following useful properties:
  • They are disjoint (see property (i) in the lemma); as explained in Section 4.2.2, having disjoint subsets that can cover a good fraction of $\mathcal{W}_{j}$ facilitates analysis.
  • While they might be unable to fully cover the set $\mathcal{W}_{j}$, these subsets taken collectively still cover a highly nontrivial fraction (i.e., at least $\frac{1}{\mathrm{poly}\log(k,T)}$) of the elements in $\mathcal{W}_{j}$.
  • Each subset $\mathcal{V}_{j}^{n}$ is linked with a suitable segment, whose associated weights have increased sufficiently from its starting point to its end point.
5.3.3 Step 3: bounding the length of segments
Recall the definition (79) of $v^{t}$ as follows
$v^{t}\coloneqq L(h^{t},w^{t})-\mathsf{OPT}\qquad\text{with }\mathsf{OPT}\coloneqq\min_{\pi\in\Delta(\mathcal{H})}\max_{1\leq i\leq k}L(h_{\pi},\mathrm{e}^{\mathsf{basis}}_{i}),$ (38)
Lemma 8.
$t_{2}-t_{1}\geq\frac{x}{2\eta}.$ (39)
Moreover, if
$\frac{qx^{2}}{50\big{(}\log_{2}(1/\eta)+1\big{)}^{2}}\geq\frac{1}{k}$ (40)
  • (a)
    the length of the segment satisfies
  • (b)
    the quantities $\{v^{t}\}$ obey
Proof.
See Appendix B.5. ∎
5.3.4 Step 4: putting all this together
To finish up, it follows from Property (ii) of Lemma 7 that
6 Limitations of proper learning?
Given that the best-known sample complexities prior to our work were derived for algorithms that either output randomized hypotheses or invoke majority votes, Awasthi et al., (2023) raised the question about how the sample complexity is impacted if only deterministic (or “proper”) hypotheses from $\mathcal{H}$ are permitted as the output of the learning algorithms. As it turns out, the restriction to proper learning algorithms substantially worsens the sample efficiency, as revealed by the following theorem.